A balloon is inflated with helium to a volume of 0.75 L at 27°C. The balloon was found several  hours later, snd the temperature had dropped to 22°C. What would be the volume of the balloon when found, if no helium had escaped? 



1
Expert's answer
2021-03-08T06:04:26-0500

V1/T1=V2/T2;

V2= V1*T2/T1=0.75*22/27=0.61 L


Answer:  the volume of the balloon is 0.61 L